Given that x is a normal variable with mean u = 47 and standard deviation = 6.2, find:
(A) P(x ≤ 120), (B) P(x ≥ 80), (C) P (108 ≤ x ≤ 117)

Answer by stanbon(75887)

Given that x is a normal variable with mean u = 47 and standard deviation = 6.2, find:
(A) P(x ≤ 120)
z(120) = (120-47)/6.2 = 11.77
P(x < 120) = P(z < 11.77) = 1
===================================
Use the same procedure on B and C.
Cheers,
Stan H.
===============

(B) P(x ≥ 80), (C) P (108 ≤ x ≤ 117)